Evaluate the following integrals.

S 5x-2 dx x2-4 s 9x+25 (x+3)2 dx 2 x3+3x2-4x-12 dx x2+x-6
Answer #1

51-2 f B 472 4 1-2 50-2 Ą x²4 (x-2)(x+2) x-2 54-2= A(+2)+6/1-2) put u=2: 10-2 = A(2+2) = A=fald A=2 put x=-2:-10-2 = B(-2-2)